OpenStudy (anonymous):

Do the three sides make a triangle ? 10, 15, 20

OpenStudy (anonymous):

I think you use the Pythagorean Theorem. \[a^2+b^2=c^2\]

OpenStudy (anonymous):

i think the formula is a^2+b^2=c^2 to find it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The sum of two sides has to be greater than the third side. sides 10, 15, and 20 10 + 15 > 20 10 + 20 >15 15 + 20 > 10 Yes, using the three sides you can make a triangle.